OpenStudy (anonymous):

Part 1: If you were solving a system of equations and you came to a statement like 4 = 4, what do you know about the system? Part 2: Solve the following system and show all of your work. 3y − 2x = 11 y + 2x = 9

OpenStudy (anonymous):

y+2x=9 -2x -2x y=9-2x replace y in the first equation for 9-2x 3(9-2x) -2x=11 27-6x-2x=11 27-8x=11 +8x +8x 27= 11+8x -11 -11 16=8x divide both sides by 8 x=2 replace x with 2 in the equation y+2x=9 y=5

OpenStudy (anonymous):

and im not sure about part 1